At present a mother is 9 times as old as her daughter. 5 Years later, the daughter's age will be 1/5th of the age of the mother. Calculate their current ages.
Answer by mananth(16946)

daughter = x years
mom=9x
after 5 years
daughter = x+5 years
Mom = 9x+5 years
1/5(9x+5)=x+5
9x+5= 5(x+5)
9x+5=4x+25
9x-4x=25-5
5x=20
/4
x=4 daughter's age
mom's age = 9*4 = 36 years
m.ananth@hotmail.ca